How do factor this polynomial completely:
14t^3 - 40t^2 + 6t^4

Thank you!

Answer by user_dude2008(1861)

14t^3 - 40t^2 + 6t^4 = 6t^4 + 14t^3 - 40t^2 = 2t^2(3t^2+7t-20) = 2t^2(t+4)(3t-5)


Answer: 14t^3 - 40t^2 + 6t^4 = 2t^2(t+4)(3t-5)